OpenStudy (anonymous):

Need help... will someone explain this to me? Solve for y. 3y – 12 = –6 A. –54 B. 18 C. 2 D. –6

OpenStudy (shubhamsrg):

Your first step should be collecting like terms together Here 3y is a different term and 12 and 6 are like terms you need 12 and 6 together on one side of the 'equal to' sign. What should you do ?

OpenStudy (anonymous):

3y – 12 = –6 So moving 12 to the other side therefore 3y = 12 -6 which is 3y = 6 and y turns out to be 2

OpenStudy (anonymous):

@Bree123

OpenStudy (anonymous):

http://openstudy.com/users/Bree123

OpenStudy (anonymous):

yes thank you